Status Report

NASA Space Station On-Orbit Status 6 Nov 2003

By SpaceRef Editor
November 6, 2003
Filed under , , ,
NASA Space Station On-Orbit Status 6 Nov 2003
iss

All ISS systems continue to function nominally, except as noted previously or below.

For CDR/SO Michael Foale, today was the first day for the HPA (Hand Posture Analysis) investigation.  After having reviewed the HPA OCBT (on-orbit computer-based training) and set up the VTR (video tape recorder) for documenting the activities, Mike used the posture acquisition glove (PAG), which has delicate sensors attached to the tops of the fingers, to operate the handgrip dynamometer (HGD).  The HGD is connected to the HPA interface box (IBOX) along with the pinch force dynamometer (PFD) and the PAG. [The research objective of the ASI/Italy (Kayser Co.)-designed HPA is to investigate the performance degradation of the human upper limb muscle-skeletal apparatus and its morphological-functional modifications during long term exposition to zero-G and to study the role of gravity in the planning and execution hierarchy of reaching, grasping, manipulating and transporting objects.  The HPA facility consists of an HGD, a PFD, and the instrumented PAG with 15 degrees of freedom, allowing the measurement of the bending angles on individual phalanxes.  PAG is coupled to a Wrist Electronic Box (WEB) housing an inertial tracking system in order to acquire tri-axial acceleration and rotation of the forearm.]

FE Alexander Kaleri continued the latest round of periodic preventive maintenance of Russian segment (RS) ventilation systems, today in the “Pirs” DC-1 docking compartment, where he cleaned the VD1 and VD2 air ducts and the V3 ventilator fan screen.  He also changed out the two PF1 & PF2 dust filters of its air ventilation system and cleaned the mesh screens of the V1 & V2 ventilator fans (last time done: 9/10).

Mike Foale recharged both batteries of the MedOps cardiac defibrillator, checked out yesterday.  Later in the day, he verified proper charge by taking voltage measurements.

At the Russian SVO-ZV water supply system, Kaleri repeated yesterday’s change-out of EDV container, hose and other equipment, and subjected the gear to another disinfection and washing cycle with water and disinfectant solution delivered on Progress-248/12P, using a hand pump and special receptacles.  The disinfection and flushing will continue tomorrow.

Today’s planned assignment (postponed from 10/29) for Mike Foale to replace RED (resistive exercise device) canisters #1009 & #1010 with cans #1001 & #1003 in the Node today was deferred again.  [Ed Lu and Yuri Malenchenko had installed the two canisters on 9/18, and #1009 had begun making scraping noise at upper load levels and intermittently at about 78% load on 10/8.]

Kaleri started another round of earth observations in the Russian Uragan (“hurricane”) program, after setting up the LIV Betacam television system at SM window #9 on his personal time.  [Today’s video targets were the Eastern coast of Central America, the Panama Canal, the Pacific Coast, landscape mapping at the Andes, and valleys of large rivers running to the East (left of track).]

Sasha also conducted the daily routine maintenance of the SM’s SOZh life support system (including ASU toilet facilities), then prepared the IMS (inventory management system) “delta” list for automatic export/import to update the database.

Mike called down the daily “ad hoc” cabin air ppO2 (oxygen partial pressure) measurements of the CSA-CP (compound specific analyzer-combustion products), used by the ground for trending analyses.

After setting up the standard PAO video equipment and preparing themselves, the crew, at 10:05am EST, participated in a live TV exchange with several hundred kids assembled for an educational event at the Space Center Houston.

The crew also received compliments from the ground on yesterday’s highly successful downlinks to three PAO events (today’s All Hands meeting of the ISS Program Office at JSC, the KSC tour of a NATO Parliamentary Assembly group on 11/10, and the fifth Anniversary of the launch of FGB “Zarya”, the first ISS element, on 11/20/1998).

At the EarthKAM payload at the Lab nadir window, Science Officer Foale changed camera lenses from 50mm to 180mm.  The changeout was timelined during a night pass to ensure that there were no pictures in the queue during the lens swap.  [EarthKAM (Earth Knowledge Acquired by Middle school students) employs an ESC 460C electronic still camera mounted at the Lab’s down-facing window, now fitted with a 180mm lens.  Powered by 16Vdc from a 120 Vdc adapter, EarthKAM takes pictures by remote operation from the ground, without crew interaction.  It is available for Middle School-ers, who submit image requests and conduct geographic research.  The requests are uplinked in a camera control file to an SSC (station support computer) laptop which then activates the camera at specified times and receives the digital images from the camera’s storage card on its hard drive, for subsequent downlink via OpsLAN.]

Both PCUs (plasma contactor units) continued in Discharge mode today, after the PHEP (Plasma Hazard Evaluation Process) Team requested a one-day extension of operations (through tomorrow).  With the recent solar upheavals, this provides researchers with additional time to assess PCU data tomorrow morning to assure that data collection was satisfactory before the PCUs are deactivated. [The two PCUs, attached to the Z1 truss (deployed on Mission 3A), emit a stream of charged xenon atoms (ions) and electrons into space.  The plasma acts as a “ground” to the environment, which prevents large charges from building up on the station, thus controlling the potential (voltage) between the ambient plasma and the ISS structure.  The current PCU operation is not done to mitigate a perceived risk, but to gain data on PCU performance for this flight attitude/configuration and the current ambient solar-event-modified ionospheric conditions.]

Both crewmembers completed their daily physical exercise program (2.5 hrs.) on TVIS treadmill, CEVIS cycle ergometer, and RED anaerobic exerciser.

Prior to providing further directions to the crew, TVIS engineers are assessing the recent crew’s calldown that the treadmill’s top wire rope loops are frayed.  [For running on the treadmill (motor-powered or passive), the crewmember wears a special harness which is hooked into strut-like SPDs (subject positioning devices), one left, one right, to keep him centered and minimize the force transferred to the station during exercise.  In addition, the runner is harnessed to the SLDs (subject load devices), which are bungees keeping his feet in contact with the running surface.]

The crewmembers again had a “free” hour each for station familiarization and acclimatization, as is standard daily rule for the first two weeks after starting station residency.

PMA-2 stowage with equipment from the Node on 11/4 was judged very successful, although the procedure took longer than the time allotted, resulting in a rescheduled PMC (private medical conference) and deferral of the SAMS ICU repair activity.  Preliminary estimates are that the stowage gain realized by the PMA-2 packing was approximately 25-27 CTB-equivalents.  Some more stowage was done today.  [Based on video and crew reports, the following items are preliminarily identified as being stowed in PMA2:  One RSP (resupply stowage platform); one BMRRM (bearing motor roll ring module) inside a bag; two triple CTBs (crew transfer bags) containing other empty CTBs; about 10 half-CTBs of towel/napkin pantry; and one EVA fluid QD tool bag (no extra tools were found in the pre-stow audit).]

TsUP/Moscow uplinked a list of stored food ration containers valid for the duration of the Increment.  [Because of the expiration dates of some food items in the Russian portion of the onboard food pantry, the uplink identified containers, which are recommended for disposal after reaching expiration.]
 
The currently ongoing solar coronal mess ejection (CME) may result in auroral displays in temperate latitudes, as far south as Florida.  Times during night passes were uplinked for the crew to observe aurorae if they are visible (opportunity lasts until sunrise on each orbit, a time duration of approximately 17 minutes). [Satellite images show that the axis of the recent auroral displays is aligned East-to-West at the latitude of Lake Michigan.]

Today’s CEO (crew earth observation) targets, taking into account the current LVLH attitude, and including the targets of the Lewis & Clark 200-year memorial locations, were Khartoum, Sudan (looking a touch right in the acute angle between the Blue and White Niles), Casablanca, Morocco (looking left midway along the smooth stretch of the Moroccan coast), Dakar, Senegal (good long oblique of the city), Albuquerque, New Mexico (looking a touch right on the Rio Grande), and Interdunal lakes, Argentina (Dynamic event.  The sunglint disc lied almost due aft and should have highlighted a mass of lakelets that pepper depressions in a major dunefield south of Buenos Aires.  Vegetated dunes with widespread lakes indicate a significant trend towards a wetter climate since the desert dunes were laid down).

CEO images can be viewed at the websites.

See also the website “Space Station Challenge” at

U.S. and Russian Segment Status (as of 1:30pm EST).

Environmental Control and Life Support (ECLSS) and Thermal Control (TCS):

  • Elektron O2 generator is powered On, 18A.  Vozdukh CO2 scrubber is On (in Manual Mode 5/3).  U.S. CDRA CO2 scrubber is on Standby (ready in dual-bed mode).  TCCS (trace contaminant control subsystem) is operating.  MCA (major constituents analyzer) is off (in Life Extending Mode).  BMP Harmful Impurities unit: absorbent bed #1 in Purify mode, bed #2 in Purify mode.  RS air conditioner SKV-1 is On; SKV-2 is Off.
  • SM Working Compartment:  Pressure (mmHg) — 742; temperature (deg C) — 27.0; ppO2 (mmHg) — 159.6; ppCO2 (mmHg) — 2.5.
  • SM Transfer Compartment:  Pressure (mmHg) — 761; temperature (deg C)  20.7.
  • FGB Cabin:  Pressure (mmHg) — 756; temperature (deg C) — 22.7.
  • Node:  Pressure (mmHg) — 748.34; temperature (deg C) — 22.9 (shell); ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • U.S. Lab:  Pressure (mmHg) — 750.07; temperature (deg C) — 25.0; ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • Joint Airlock (Equip. Lock):   Pressure (mmHg) — 751.17; temperature (deg C) — 23.9; shell heater temp (deg C) — 22.0, ppO2 (mmHg) — n/a; ppCO2 (mmHg) — n/a.
  • PMA-1:  Shell heater temp (deg C) — 22.9
  • PMA-2:  Shell heater temp (deg C) — 20.6

(n/a = data not available)

Electrical Power Systems (EPS):

  • Both P6 channels fully operational.  BGA (beta gimbal assembly) 2B and 4B both in Directed (dual-angle) Position (non-suntracking). 
  • SM batteries:   Battery #8, formerly known as #7, is off (still in slot #8); battery #6 is in “Cycle” mode; all other batteries (6) are in “Partial Charge” mode (batteries #1 and #3 are degraded). 
  • FGB batteries:   Batteries #1 is off; all other batteries (5) are in “Partial Charge” mode.
  • Plasma Contactor Unit PCU-1 in Discharge mode; PCU-2 in Discharge mode.

Command & Data Handling Systems:

  • C&C-2 MDM is prime, C&C-1 is back-up, and C&C-3 is in standby.
  • GNC-1 MDM is prime; GNC-2 is Backup.
  • INT-2 is operating; INT-1 is Off.
  • EXT-2 is On (primary), EXT-1 is Off (both now upgraded to R3).
  • LA-1, LA-2 and LA-3 MDMs are all operating.
  • PL-1 MDM is Off; PL-2 MDM is Operational.
  • APS-1 (automated payload switch #1) and APS-2 are both On.
  • SM Terminal Computer (TVM): 3 redundant lanes (of 3) operational.
  • SM Central Computer (TsVM): 2 redundant lanes (of 3) operational; string #3 dropped out 10/22.

Propulsion System:

  • Total propellant load available: 3856 kg (8501b) as of 10/23  [SM(755) + FGB(2749) + Progress M(352) + Progress M-1(0)].  (Capability: SM — 860 kg; FGB — 6120 kg).

Attitude Control Systems:

  • 3 CMGs on-line (CMG-1 failed).
  • State vector source — U.S. SIGI-1 (GPS)
  • Attitude source — U.S. SIGI-1 (GPS)
  • Angular rate source — RGA-1

Flight Attitude:

  • LVLH YVV (y-axis in velocity vector, i.e. flying “sidewise”) [yaw: -90 deg, pitch: -8.9 deg., roll: 1.8 deg]), with CMG Momentum Management.  LVLH until 11/8.

Communications & Tracking Systems:

  • FGB MDM-1 is powered Off; FGB MDM-2 is operational.
  • All other Russian communications & tracking systems are nominal.
  • S-band is operating nominally (on string 2).
  • Ku-band is operating nominally.
  • Audio subsystem is operating nominally (IAC-2 is prime, IAC-1 is suspect).
  • Video subsystem operating nominally.
  • HCOR (high-rate communications outage recorder) is operating nominally.

Robotics:

  • SSRMS/Canadarm2 based at MBS PDGF #2/LEE A, with Keep Alive (KA) power on both strings.
  • MBS: KA power on both strings. 
  • MT: latched and mated at WS4. 
  • POA: KA power on both strings.
  • RWS (robotics workstations): Lab RWS is On (DCP connected); Cupola RWS is Off.

ISS Orbit  (as of yesterday afternoon, 2:16pm EST [= epoch]):

  • Mean altitude — 376.9 km
  • Apogee — 381.4 km
  • Perigee — 372.5 km
  • Period — 92.1 min.
  • Inclination (to Equator) —  51.63 deg
  • Eccentricity — 0.0006649
  • Orbits per 24-hr. day — 15.64
  • Mean altitude loss in last 24 hours — 200 m
  • Revolutions since FGB/Zarya launch (Nov. ’98)  — 28313
  • For more on ISS orbit and worldwide naked-eye visibility dates/times, see http://www.hq.nasa.gov/osf/station/viewing/issvis.html

SpaceRef staff editor.